OpenStudy (anonymous):

how do you solve this problem: parallel to y=2x+5 and passes through (-1, 3)

OpenStudy (immanuelv):

switch x and y first x=2y+5 then solve for y 2y=5-x then y=(5-x)/2 so answer is (5-x)/2

OpenStudy (mertsj):

Parallel lines have equal slopes. The slope of y = 2x+5 is 2 so write the equation of a line whose slope is 2 and contains the point (-1,3) y-3=2(x+1) y - 3 = 2x+2 y = 2x+5